PLEASE SHOW WORK ON HOW TO FIND P-VALUE

17. In a study of memory recall, 5 people were given 10 minutes to memorize a list of 20 words. Each was asked to list as many of the words as he or she could remember both 1 hour and 24 hours later. The result is shown in the following table.
Number of Words Recalled
Subject
1 hour later
24 hours later
1
14
12
2
18
15
3
11
9
4
13
12
5
12
12
Is there evidence to suggest that the mean number of words recalled after 1 hour exceeds the mean recall after 24 hours?
Assume we want to use a 0.05 significance level to test the claim.

(e) Determine the p-value. Round your answer to three decimal places. Show all work; writing the correct critical value, without supporting work, will receive no credit
Answer #1

Ans:

1 hr 24 hr d
1 14 12 2
2 18 15 3
3 11 9 2
4 13 12 1
5 12 12 0
mean(d)= 1.6
std. dev.= 1.140

Test statistic:

t=(1.6-0)/(1.14/sqrt(5))

t=3.138

df=5-1=4

p-value=tdist(3.138,4,1)=0.0175

As,p-value<0.05,we reject the null hypothesis.

There is sufficient evidence to conclude that  the mean number of words recalled after 1 hour exceeds the mean recall after 24 hours.